OpenStudy (anonymous):

$14.99 clock; 10% discount = ?

OpenStudy (venomblast):

multiply then subtract from 14.99

OpenStudy (anonymous):

HELP!!!!! =0

jimthompson5910 (jim_thompson5910):

what is 10% of $14.99 ?

OpenStudy (venomblast):

remember 10%= .1 multiply that from 14.99 after getting the answer, subtract it from 14.99

OpenStudy (mathstudent55):

10% discount means you pay the original price minus 10% of the original price. First, find 10% of $14.99 Then subtract that from $14.99

OpenStudy (venomblast):

remember to use the money sign($)

OpenStudy (anonymous):

kk

OpenStudy (mathstudent55):

10% of $14.99 "%" means "divide by 100" That means that 10% means 10/100 = 1/10 = 0.1 "of" in math usually means "times" That means that 10% of $14.99 means 0.1 * $14.99

OpenStudy (mathstudent55):

To find the discount, multiply 0.1 * 14.99 Then subtract that amount from 14.99 to find the discounted price.

OpenStudy (anonymous):

got it ;)

OpenStudy (anonymous):

so about 13.50

OpenStudy (mathstudent55):

Yes, it's $13.49

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

thanks : )